OpenStudy (anonymous):

R=1/2(r1+r2) solve for r2

OpenStudy (anonymous):

Congrats on your first question! :) There are a few things for this problem you need to do. Are you ready?

OpenStudy (anonymous):

1/r=1/r1+1/r2 Get rid of the denominators; multiply thru by (R*R1*R2), resulting in: R1*R2 = R*R2 + R*R1 : Factor out R on the right: R1*R2 = R(R2 + R1) : Divide both sides by (R2 + R1) r1 times r2/r1+ r2 = R : The old "product over sum" formula for two resistors in parallel, right?

OpenStudy (anonymous):

Did you follow that?

OpenStudy (anonymous):

@Math.... Math is really fun once you understand it!